OpenStudy (anonymous):

The sides of a square are five to the power of two fifths inches long. What is the area of the square? A. five to the power of four fifths square inches B. five to the power of four twenty fifths square inches C. twenty five to the power of four fifths square inches D. twenty five to the power of four twenty fifths square inches

OpenStudy (anonymous):

To calculate the area of the square, multiply\[5^{\frac{ 2 }{ 5 }} \times 5^{\frac{ 2 }{ 5 }}\]What do you get?

OpenStudy (triciaal):

when you multiply numbers/variables with the same base add the exponents

OpenStudy (anonymous):

Sorry was away for 14 hours I got the question, Thank's to both of you!